OpenStudy (anonymous):

Catherine ate 1/6 of cake and gave her friends 2/5 of the remainder. what fraction of cake was left?

OpenStudy (anonymous):

1/2

ganeshie8 (ganeshie8):

2/3

OpenStudy (anonymous):

17/30

OpenStudy (anonymous):

i would show you how if I knew \( \LaTeX) Aww mann... Not again....

Parth (parthkohli):

^ \(\LaTeX\)

OpenStudy (anonymous):

making me cry, Parth ;(

OpenStudy (anonymous):

13/30

OpenStudy (anonymous):

2/5 of the remainder, parth.

ganeshie8 (ganeshie8):

1 - (1-1/6)x2/5 @ParthKohli wrong expression check again

Parth (parthkohli):

i see, we'll multiply as 'of' means multiply.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large (1 - {1 \over 6}) \times 0.4 }\)

OpenStudy (anonymous):

can u please explain

OpenStudy (anonymous):

+1/6 to your answer, @ganeshie8

OpenStudy (anonymous):

you didnt count cat's share.

OpenStudy (anonymous):

sorry dont follow.. and you show me your solution? plz ^^

OpenStudy (anonymous):

wait, i'll try.

OpenStudy (anonymous):

Let the whole cake is of x size... So, Catherine ate 1/6 of x = x/6 Cake remains = x - x/6 = 5x/6.. Friends ate = 2/5 of Remainder = 2/5 of 5x/6 = x/3.. So, Cake left = x - x/6 - x/3 = (6x - x - 2x)/6 = 3x/6 = x/2.. So, 1/2 of Cake is left behind....

OpenStudy (anonymous):

cat's friends get (2/5)of the remaining =>(2/5)(1 - 1/6) =>1/3 of the cake. so the fraction of the cake that has vanished=1/3 + 1/6 =1/2

OpenStudy (anonymous):

ohhhhhhhhhhh TY ^^ know i know